OpenStudy (anonymous):

divide and simplify √6x^2/√25x^4

OpenStudy (e.cociuba):

Answerer 1 √6x^2 / √25x^4 = x√6/5x^2 = √6/5x Answerer 2 sqrt(6)/5*x